OpenStudy (anonymous):

Three times a first number is decrease by second number is 22. The first number is increase by twice number is -9. Find the two numbers

OpenStudy (anonymous):

3x-y=22 x+2y=-9

OpenStudy (anonymous):

solve the system

OpenStudy (anonymous):

try by multiplying 1º equation by 2 and later adding bouth equations

OpenStudy (anonymous):

I got 5,-7

OpenStudy (anonymous):

6x-2y=44 x+2y =-9 ----------- + 7x=35 so x=5 from 2º equation: 5+2y=-9 2y=-14 so y =-7 you are right